# Balanced Diffusion-Guided Fusion for Multimodal Remote Sensing Classification

### v1

@@ IV-E Analysis of the Computational Complexity
In this section, we evaluate the computational complexity of all models in terms of giga floating-point operations (GFLOPs) and number of parameters (millions). Table VII presents these metrics for the four datasets considered. Note that SS-MAE and BDGF employ a pre-training paradigm, and (P) and (T) denote pre-training and training phases, respectively. Under unsupervised learning, larger models generally generalize better [55, 56]. Pre-training methods incur higher computational costs than other approaches, and adopting a dual-branch structure for the pre-training autoencoder and diffusion model further increases complexity. SS-MAE and BDGF mitigate this issue by merging multimodal images along the channel dimension. Their superior classification results across the four datasets confirm the effectiveness of pre-training. For the proposed BDGF, the parameter count arises from integrating multi-branch sub-networks, while the FLOPs reflect the diffusion model’s pre-training. Although BDGF exhibits higher computational complexity, it remains within acceptable limits and achieves the best classification performance.
IV-F Analysis of the Transferability
Different from the fine-tuning training paradigm such as SS-MAE, the proposed BDGF framework employs an unsupervised diffusion process to learn the joint data distribution of multimodal images and directly leverages these learned features for downstream classification. To evaluate both the transferability of our adaptive masking strategy and the modularity of our multi-branch network, we perform cross-model feature-transfer experiments with SpectralDiff, which shares a similar training paradigm. Specifically, we first adapt SpectralDiff to multimodal inputs using early fusion via channel-wise concatenation [9, 18, 43]. From the pre-trained diffusion backbones of both methods, we extract features and graft them into the other model’s classification head. In Table VIII, the pre-trained diffusion model and sub-network of SpectralDiff are denoted as “Concatenate Diff” and “Attention Classified Module”, respectively, while the corresponding annotations of our method are “Masking Diff” and “Multi-Branch Network”.

### v2

@@ IV-F Analysis of the Transferability
Different from the fine-tuning training paradigm such as SS-MAE, the proposed BDGF framework employs an unsupervised diffusion process to learn the joint data distribution of multimodal images and directly leverages these learned features for downstream classification. To evaluate both the transferability of our adaptive masking strategy and the modularity of our multi-branch network, we perform cross-model feature-transfer experiments with SpectralDiff [32], which shares a similar training paradigm. Specifically, we first adapt SpectralDiff to multimodal inputs using early fusion via channel-wise concatenation [8, 16, 44]. From the pre-trained diffusion backbones of both methods, we extract features and graft them into the other model’s classification head. In Table VIII, the pre-trained diffusion model and sub-network of SpectralDiff are denoted as “Concatenate Diff” and “Attention Classified Module”, respectively, while the corresponding annotations of our method are “Masking Diff” and “Multi-Branch Network”.